OpenStudy (anonymous):

\[f(x)=4(x+1)^{2}-2\] vertex?

OpenStudy (amistre64):

yeah... -1,-2

myininaya (myininaya):

(-1,-2)

OpenStudy (anonymous):

x-intercept of parabola?

myininaya (myininaya):

f(x)=a(x-h)^2+k gives vertex(h,k) f(x)=a(x+h)^2+k gives vertex(-h,k) f(x)=a(x-h)^2-k gives vertex (h,-k) f(x)=a(x+h)^2-k gives vertex (-h,-k)

myininaya (myininaya):

to find the x-intercept, set y=0 and solve for x

OpenStudy (amistre64):

y = 0 when 1/2 = (x+1)^2 -1 +- sqrt(1/2)